V. Kokilashvili

A. Razmadze Mathematical Institute, Georgian Academy of Sciences Tbilisi, Georgia

A SURVEY OF RECENT RESULTS ON THE BOUNDARY VALUE PROBLEMS FOR HOLOMORPHIC FUNCTIONS

In our lecture we give a survey of the results obtained for the recent years by the participants of seminars held at A. Razmadze Mathematical Institute on boundary value problems for holomorphic and harmonic functions and on singular integral equations. The original source of these investigations turned out to be the richest scientific hertigate of N. Muskhelishvili.